OpenStudy (anonymous):

Choose the correct answer. Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores? A. Joelle: 140, Anita: 240 B. Joelle: 180, Anita: 245 C. Joelle: 270, Anita: 155 D. Joelle: 155, Anita: 270

OpenStudy (anonymous):

I think its D

OpenStudy (nurali):

Let x = Anita's score and y = Joelle's score then x + y = 425 (equation 1) x = 2y - 40 (equation 2) Use the definition of x from equation 2 and substitute it into equation 1: x + y = 425 2y - 40 + y = 425 3y - 40 = 425 3y = 465 y = 155 (Joelle's score) . Anita's score: x + y = 425 x + 155 = 425 x = 270 (Anita's score)

OpenStudy (nurali):

Ur answer is correct.